# Gibbs polystability of Fano manifolds, stability thresholds and symmetry breaking

### v1

@@ 1.7. Organization of the paper
  • In Section 4 applications of Theorem 3.13 to logarithmic Hardy–Littlewood–Sobolev inequalities and Moser–Trudinger inequalities are given and a phenomenon of spontaneous symmetric breaking is exhibited.
@@ 2.2.3. The Fulton–MacPherson compactification of configuration spaces
where $\pi_{W}\colon\mathrm{Bl}_{W}(X^{N})\rightarrow X^{N}$ is the blow-up morphism, whose exceptional divisor we denote by $E_{\pi_{W}}.$ It will thus be enough to show that $f_{*}^{-1}(E_{\pi_{W}})=E_{W^{\prime}},$ where $f_{*}^{-1}(E_{\pi_{W}})$ denotes the strict transform of $E_{\pi_{W}}$ under $f.$ To this end let us first observe that there exists $W^{\prime}$ in $\mathcal{W}$ such that
@@ 2.3.3. The case of non-trivial $\mathcal{G}$
$\min_{W\in\mathcal{W}_{p}^{\mathrm{ss}}}\frac{\mathrm{codim}W-w\mathrm{codim}W}{\frac{1}{k}\mathrm{codim}W(\mathrm{codim}W-1)/2}$ $=\min_{W\in\mathcal{W}_{p}^{\mathrm{ss}}}\frac{2k(1-w)}{\mathrm{codim}W-1}$
$=\frac{2(N-1)(1-w)}{V(\lfloor N/2\rfloor-1)}=\frac{N-1}{(\lfloor N/2\rfloor-1)}.$
@@ 3.1. Setup
Furthermore, we fix a $\mathcal{K}$–invariant metric $\phi_{0}$ on $K_{(X,\Delta)}$ with positive curvature form $\omega_{0}(=dd^{c}\phi_{0})$ (using additive notation for metrics on line bundles, recalled below). The fixed metric $\phi_{0}$ on $K_{(X,\Delta)}$ corresponds to a volume form on $X$ with divisorial singularities that we denote by $\mu_{0}$ (see Section 3.2.1). We will denote by $\mathcal{P}(X)$ the space of all probability measures on $X,$ endowed with its weak topology and by $\mathcal{P}(X)_{0}$ the subspace of all measures $\mu$ with vanishing moment:
@@ 3.2.3. Pluripotential theory and energy
In this section we recall the notion of pluricomplex energy of a probability measure $\mu$ introduced in [23], but using the notation in [12, 13]. Given a Kähler form $\omega_{0}$ on a compact complex manifold consider the space of all Kähler potentials:
and denote by $\text{PSH}(X,\omega)$ its closure in $L^{1}(X).$ Given a line bundle $L$ over $X$ and a metric $\phi_{0}$ on $L$ with curvature form $\omega_{0}$, expressing $u=\phi-\phi_{0}$ for a metric $\phi$ on $L,$ we may identify the spaces $\mathcal{H}(X,\omega_{0})$ and $\text{PSH}(X,\omega_{0})$ with $\mathcal{H}(L)$ and $\text{PSH}(\mathcal{L}),$ respectively. We also set
Thus, $\mathcal{H}(X,\omega_{0})\subset\mathcal{H}_{1,1}(X,\omega)\subset\text{PSH}(X,\omega_{0})$. Given a log Fano manifold $(X,\Delta)$ we will take $L=-K_{(X,\Delta)}$ and $\phi_{0}$ as in the setup introduced above. Consider the following functional on $C^{\infty}(X)$:
@@
As a consequence, the restriction of $\mathcal{E}$ to $\mathcal{H}(X,\omega_{0})$ is increasing and thus admits a canonical extension to $\text{PSH}(X,\omega_{0}),$ namely the smallest extension which is upper semi-continuous with respect to the $L^{1}$–topology [23]. The subspace of $\text{PSH}(X,\omega_{0})$ consisting of all $u$ such that $\mathcal{E}(u)>$-$\infty$ is denoted by $\mathcal{E}^{1}(X,\omega_{0}).$ Formula3.4 still applies on $\mathcal{E}^{1}(X,\omega_{0}),$ when $\omega_{u}^{n}$ is the Monge–Ampère measure defined in terms of non-pluripolar products [23].
@@
For example, any measure $\mu$ on $X$ with divisorial singularities (or more, generally with an $L^{p}$–density for some $p>1)$ has finite energy, $E(\mu)<\infty.$ Thus the properties 3.8 and 3.9 determine $E$ on $\mathcal{P}(X)\cap L_{\text{loc}}^{p}$ for any $p>1$ (and the functional $E$ on $\mathcal{P}(X)$ can then be characterized as the greatest lower semi-continuous extension of $E$ from $\mathcal{P}(X)\cap L_{\text{loc}}^{p}$ to $\mathcal{P}(X)$).
@@
$V\mathcal{E}(u)=-\frac{1}{2}\int_{X}du\wedge d^{c}u+\int_{X}u\omega_{0},\,\,2VE(\mu)=-\int_{X}u_{\mu}\mu=\int_{X}du_{\mu}\wedge d^{c}u_{\mu},$
if $E(\mu)<\infty$ or equivalently $u_{\mu}$ is in the space $H^{1}(X)\subset L^{2}(X)$ of all $u$ such that $du\in L^{2}(X).$ Moreover, in this case the sup in formula 3.5 can be taken over all of $C^{0}(X)$ or over $H^{1}(X).$ Indeed, by classical Hilbert space theory the latter sup is attained by $u_{\mu}\in H^{1}(X)$ solving the (Laplace) equation 3.6 and hence $u_{\mu}\in\text{PSH}(X,\omega_{0})$.
@@ 3.2.4. The space $\mathcal{R}_{1,1}(X,\omega)$ of geodesics rays in $\mathcal{H}_{1,1}(X,\omega)$
The space $\mathcal{H}(X,\omega_{0})$ can be endowed with a canonical Riemannian metric, first introduced by Mabuchi, where the squared length of a tangent vector $\dot{u}$ at $u$ is defined by $\int_{X}\dot{u}^{2}\omega_{u}^{n}.$ The corresponding notion of geodesics was extended to $\mathcal{H}_{1,1}(X,\omega_{0})$ in [40]. We will denote by $\mathcal{R}(X,\omega)_{1,1}$ the corresponding space of all geodesic rays $u_{t}$ in $\mathcal{H}_{1,1}(X,\omega),$ parametrized by $t\in[0,\infty[.$
3.2.5. The Mabuchi functional $\mathcal{M}$ versus the free energy functional $F$
@@
We recall that the function $\mu\mapsto D(\mu|\nu)$ is lower semicontinuous and convex on $\mathcal{P}(X)$ as follows, for example, from the following well-known Legendre duality formula:
@@ 3.2.6. The Futaki character of $(X,\Delta)$ and its quantization
More generally, 3.17 holds if $\psi$ and $\phi$ have finite energy and 3.18 holds if $(dd^{c}\psi)^{n}/V$ has finite entropy. In particular, if $(X,\Delta)$ admits a KE metric, then $\text{Fut}_{(X,\Delta)}(\cdot)\equiv 0.$
@@
This is well-known, at least for smooth data, but for completeness we provide simple proofs. To prove 3.17 first note that, for any given any $g\in\mathcal{G},$
@@
This follows directly from the fact that the differential $d\mathcal{E},$ which is the one-form on 3.4 defined by formula 3.4, satisfies $g^{*}(d\mathcal{E})=d(g^{*}\mathcal{E}).$ Replacing $g$ with a one-parameter subgroup $g_{\tau}$ and differentiating with respect to $\tau$ then proves formula 3.17. Next, observe that $\text{Fut}_{(X,\Delta)}(V^{1,0})=0$ for all $V^{1,0}\in\mathfrak{g}$ iff $\int_{X}\mathcal{L}_{\text{Re}V}\phi_{0}(dd^{c}\phi_{0})^{n}=0$ for all $V^{1,0}\in\mathfrak{g}$ (by applying $\text{Fut}_{(X,\Delta)}$ to both $V^{1,0}$ and $iV^{1,0}).$ Now, using that $\mathfrak{g}$ is reductive, we can decompose $\text{Re}V=W_{1}+JW_{2}$ where $W_{i}\in\mathfrak{k}$, the Lie algebra of $\mathcal{K}$. Since $\phi_{0}$ is assumed $\mathcal{K}$–invariant $\text{Fut}_{(X,\Delta)}(V^{1,0})=\int h_{W_{2}}(\omega_{0})^{n}/V,$ which shows that $\text{Fut}_{(X,\Delta)}(\cdot)\equiv 0$ iff $\int_{X}\boldsymbol{m}(\omega_{0})^{n}/V=0,$ i.e. iff $\omega_{0}^{n}/V\in\mathcal{P}(X)_{0}.$ Finally, applying formula 3.11, we can decompose
@@
Differentiating with respect to $t$ and using 3.3 thus proves 3.18. More generally, the proof of 3.17 only used that $\phi$ and $\psi$ have finite energy. Moreover, when $(dd^{c}\psi)^{n}$ has finite entropy it is shown in [23] that $\psi$ has finite energy, showing that $D(dd^{c}\psi)^{n}|\mu_{\psi})<\infty.$ Finally, if $(X,\Delta)$ admits a Kähler–Einstein metric $\psi,$ then, by [24], $(dd^{c}\psi)^{n}$ has finite entropy and $\psi$ minimizes the Mabuchi functional on the space of metrics with finite energy. Hence, $\frac{d}{dt}\mathcal{M}\left(\exp(t\text{Re}V^{1,0})^{*}\psi\right)_{|t=0}$, showing that $\text{Fut}_{(X,\Delta)}(\cdot)\equiv 0.$
@@ 3.2.7. The moment maps $\boldsymbol{m}$ and $\boldsymbol{m}_{N}$
where $h_{W}$ denotes the Hamiltonian of a real vector field $W$ on $X$ corresponding to an element $W$ in $\mathfrak{k},$ defined by $h_{W}\coloneqq\mathcal{L}_{JW}\phi$ (defined as in formula 3.16). We will denote by $\boldsymbol{m}_{N}$ the moment map of the diagonal action of the group $\mathcal{K}$ on $X^{N},$ endowed with the symplectic form
@@ 3.3. The structure of the space $\mathcal{P}(X)_{0}$ of probability measures with vanishing moment
and since $i\partial\bar{\partial}\phi^{\tau}$ is smooth and strictly positive on $X-\Delta$ it follows that $W=0$ on $X-\Delta.$ Thus, by continuity, $W=0$ on $X,$ which is a contradiction. Hence, 3.21 holds. Finally, the inequality 3.22 is shown in essentially the same way as in the proof of [21, Lemma 9.1]. More precisely, in [21, Lemma 9.1] it is assumed that the corresponding metric $\phi^{\tau}$ is independent of the real part of $\tau$ (in our notation) and that $\phi^{\tau}$ satisfies the complex geodesic equation, that is the local function $\Phi(z,\tau)\coloneqq\phi^{\tau}(z)$ satisfies
@@
where $h_{W}$ is the Hamiltonian function corresponding to $W$ and $\omega_{0}.$ First recall that $h_{W}\coloneqq\mathcal{L}_{JW}\phi_{0},$ (see Section 3.2.7). To prove formula 3.24 note that, by 3.9, denoting by $\phi$ the potential of $\mu,$
@@
which proves 3.24, using the previous lemma and the assumption that the Futaki character of $(X,\Delta)$ vanishes.
@@
First note that $F_{-\gamma}$ is $\mathcal{K}$–invariant, since $\phi_{0}$ is $\mathcal{K}$–invariant. Moreover, by definition, $F_{-\gamma}=(1-\gamma)E+F_{-1}.$ Since $F_{-1}$ is $\mathcal{G}$–invariant, under the assumption about the vanishing of the Futaki character, the first statement of the lemma thus follows and, as a consequence, also the first identity in formula 3.25. Thus, to prove the rest of the identities 3.25, it will be enough to show that for any given $\mu$ in $\mathcal{P}(X)$ such that $F_{\gamma}(\mu)<\infty$, there exists $\nu_{j}\in\mathcal{V}(X)$ such that $F(\nu_{j})\rightarrow F(\mu).$ But this follows from [27, Lemma 3.1], implying that for any given $\mu$ such that $E(\mu)<\infty$ and $D(\mu|\mu_{0})<\infty,$ there exists a sequence $\mu_{j}$ converging towards $\mu$ in $\mathcal{P}(X)$ such that $E(\mu_{j})$ and $D(\mu_{j}|\mu_{0})$ converge towards $E(\mu)$ and $D(\mu|\mu_{0}),$ respectively. ∎
@@
To prove the first statement of the proposition it is enough to show that when changing the metric $\phi_{0}$ the corresponding pluricomplex energy $E_{\phi_{0}}$ only changes by an additive bounded term. To this end note that, by formula 3.9, given $\phi_{0}$ and $\psi_{0}$ in $\mathcal{H}(L)$ the differential $d(E_{\phi_{0}}-E_{\phi_{1}})|_{\mu}$ is represented by the function $u\coloneqq\phi_{0}-\phi_{1},$ for any $\mu$ in $\mathcal{P}(X)$ satisfying $E(\mu)<\infty.$ Hence, there exists a constant $C$ such that
@@
showing that $\left|E_{\phi_{0}}(\mu)-E_{\phi_{1}}(\mu)\right|\leq C+\left\|u\right\|_{L^{\infty}},$ as desired. Finally, the second statement follows from combining formulae 3.11 and 3.10 with proposition 3.4 and using that $F_{-1}$ is $\mathcal{G}$–invariant (Section 3.2.6). ∎
@@
for some positive constants $A$ and $B.$ Indeed, by 3.10, $E(\frac{\omega_{u}^{n}}{V})=(\mathcal{I}-\mathcal{J})(u)$, which is comparable to $\mathcal{J}(u)$ [10, Lemma 2.13] and by [48, Prop 5.5] the inequalities 3.28 do hold when $E(\frac{\omega_{u}^{n}}{V})$ is replaced by $\mathcal{J}(u).$ This proves $1\iff 2$.
Next, note that, $3\implies 2$ follows from the previous proposition. To prove that $2\implies 3$ we recall that it was pointed out in [48, Remark 3.9] that, by density, the inequality 3.27 is equivalent to the corresponding inequality on the metric completion of the metric space $\left(\mathcal{H}(X,\omega_{0}),d_{1}\right),$ which, by [46], may be identified with $\mathcal{E}^{1}(X,\omega_{0}).$ Combined with the inequalities 3.28 this shows that (after perhaps decreasing $\epsilon$ and increasing $C$) there exist $\epsilon,C>0$ such that
@@
To prove the last statement of the theorem, recall that any KE volume form $\mu$ in $\mathcal{P}(X)$ minimizes $F_{-1}$ on $\mathcal{P}(X)$ [24]. Thus, by Lemma 3.5, there exists a KE volume form $\pi_{0}(\mu)$ minimizing $F_{-1}$ on $\mathcal{P}(X)_{0}$ (but, a priori, $\pi_{0}(\mu)$ is only determined modulo $\mathcal{K}).$ To prove that there exists a unique KE volume form $\mu_{\text{KE}}$ in $\mathcal{P}(X)_{0},$ we recall that, as shown in [21, Lemma 6.3], for any given $\phi_{0}\in\mathcal{H}(-K_{(X,\Delta)})$ there exists a unique KE volume form $\mu_{-1}$ in $\mathcal{P}(X)$ minimizing $E$ on the space of KE volume forms in $\mathcal{P}(X),$ i.e. on $\mathcal{G}\mu_{-1}.$ Moreover, $\mu_{-1}$ arises as the limit 3.26. In the present case, where $\phi_{0}$ is assumed $\mathcal{K}$–invariant, it thus follows from Prop 3.4 that $\mu_{-1}\in\mathcal{P}(X)_{0}$ and conversely, any KE volume form in $\mathcal{P}(X)_{0}$ coincides with $\mu_{-1}.$ It also follows from the aforementioned uniqueness that $\mu_{\text{KE}}$ is $\mathcal{K}$–invariant. ∎

### v2

@@ 1.7. Organization of the paper
  • In Section 4 applications of Theorem 3.13 to constrained logarithmic Hardy–Littlewood–Sobolev inequalities and Moser–Trudinger inequalities are given and a phenomenon of spontaneous symmetric breaking is exhibited.
  • In Section 5 we establish a sharp version of the constrained logarithmic Hardy–Littlewood–Sobolev inequalities on the two-sphere, by combining Theorem 3.13 with the uniqueness result for the Liouville equation on $S^{2}$ established in [72].
@@ 2.2.3. The Fulton–MacPherson compactification of configuration spaces
where $\pi_{W}\colon\mathrm{Bl}_{W}(X^{N})\rightarrow X^{N}$ is the blow-up morphism, whose exceptional divisor we denote by $E_{\pi_{W}}.$ It will thus be enough to show that $f_{*}^{-1}(E_{\pi_{W}})=E_{W},$ where $f_{*}^{-1}(E_{\pi_{W}})$ denotes the strict transform of $E_{\pi_{W}}$ under $f.$ To this end let us first observe that there exists $W^{\prime}$ in $\mathcal{W}$ such that
@@ 2.3.3. The case of non-trivial $\mathcal{G}$
$\min_{W\in\mathcal{W}_{p}^{\mathrm{ss}}}\frac{\mathrm{codim}W-w\mathrm{codim}W}{\frac{1}{k}\mathrm{codim}W(\mathrm{codim}W-1)/2}=\min_{W\in\mathcal{W}_{p}^{\mathrm{ss}}}\frac{2k(1-w)}{\mathrm{codim}W-1}=\frac{2(N-1)(1-w)}{(\lfloor N/2\rfloor-1)V}=\frac{N-1}{(\lfloor N/2\rfloor-1)}$
@@ 3.1. Setup
Furthermore, we fix a $\mathcal{K}$–invariant metric $\phi_{0}$ on $K_{(X,\Delta)}$ with positive curvature form $\omega_{0}(=dd^{c}\phi_{0})$ (using additive notation for metrics on line bundles, recalled below). The fixed metric $\phi_{0}$ on $K_{(X,\Delta)}$ corresponds to a volume form on $X$ with divisorial singularities that we denote by $\mu_{0},$ which (up to scaling the metric on $K_{(X,\Delta)}$)) we may assume is a probability measure; see Section 3.2.1. We will denote by $\mathcal{P}(X)$ the space of all probability measures on $X,$ endowed with its weak topology and by $\mathcal{P}(X)_{0}$ the subspace of all measures $\mu$ with vanishing moment:
@@ 3.2.2. Kähler–Einstein metrics for $(X,\Delta)$
3.2.3. Kähler metrics vs Kähler potentials and metrics on line bundles
Given a Kähler form $\omega_{0}$ on a compact complex manifold, consider the space of all Kähler potentials:
Using the map $u\mapsto\omega_{u}$ we will identify the space of all Kähler metrics $\omega$ in the cohomology class $\omega_{0}]\in H^{2}(X,\mathbb{R})$ with $\mathcal{H}(X,\omega_{0})/\mathbb{R},$ under the additive action of $\mathbb{R}$ on $\mathcal{H}(X,\omega_{0})$ (acting as $u\rightarrow u+C).$ Moreover, given a line bundle $L$ over $X$ and a metric $\phi_{0}$ on $L$ with curvature form $\omega_{0}$, expressing
(3.3) $u=\phi-\phi_{0}$
for a metric $\phi$ on $L,$ we may identify the space $\mathcal{H}(X,\omega_{0})$ with the space with $\mathcal{H}(L)$ of smooth metrics on $L$ with positive curvature form.
@@ 3.2.4. Pluripotential theory and energy
In this section we recall the notion of pluricomplex energy of a probability measure $\mu$ introduced in [23], but using the notation in [12, 13]. Given a Kähler form $\omega_{0}$ on a compact complex manifold $X,$ denote by $\text{PSH}(X,\omega)$ the closure of $\mathcal{H}(X,\omega_{0})$ in $L^{1}(X),$ called the space of all $\omega_{0}$-plurisubharmonic functions. Given a line bundle $L$ over $X$ and a metric $\phi_{0}$ on $L$ with curvature form $\omega_{0}$ we can identify $\text{PSH}(X,\omega_{0})$ with $\mathcal{H}(L)$ and $\text{PSH}(\mathcal{L}),$ using formula 3.3. Given a log Fano manifold $(X,\Delta)$ we will take $L=-K_{(X,\Delta)}$ and $\phi_{0}$ as in the setup introduced above. Consider the following functional on $C^{\infty}(X)$:
@@
As a consequence, the restriction of $\mathcal{E}$ to $\mathcal{H}(X,\omega_{0})$ is increasing and thus admits a canonical extension to $\text{PSH}(X,\omega_{0}),$ namely the smallest extension which is upper semi-continuous with respect to the $L^{1}$–topology [23]. The subspace of $\text{PSH}(X,\omega_{0})$ consisting of all $u$ such that $\mathcal{E}(u)>$-$\infty$ is denoted by $\mathcal{E}^{1}(X,\omega_{0}).$ Formula3.5 still applies on $\mathcal{E}^{1}(X,\omega_{0}),$ when $\omega_{u}^{n}$ is the Monge–Ampère measure defined in terms of non-pluripolar products [23].
@@
For example, any measure $\mu$ on $X$ with divisorial singularities (or more, generally with an $L^{p}$–density for some $p>1)$ has finite energy, $E(\mu)<\infty.$ Thus the properties 3.9 and 3.10 determine $E$ on $\mathcal{P}(X)\cap L_{\text{loc}}^{p}$ for any $p>1$ (and the functional $E$ on $\mathcal{P}(X)$ can then be characterized as the greatest lower semi-continuous extension of $E$ from $\mathcal{P}(X)\cap L_{\text{loc}}^{p}$ to $\mathcal{P}(X)$).
@@
(3.12) $\mathcal{E}(u)=-\frac{1}{2V}\int_{X}du\wedge d^{c}u+\int_{X}u\frac{\omega_{0}}{V},\,\,E(\mu)=\frac{1}{2V}\int_{X}du_{\mu}\wedge d^{c}u_{\mu}=-\frac{1}{2}\int_{X}u_{\mu}\mu$
if $E(\mu)<\infty$ or equivalently $u_{\mu}$ is in the Sobolev space $W^{2,1}(X)$ of all $u\in L^{2}(X)$ such that $du\in L^{2}(X).$ Moreover, when $n=1$ the sup in formula 3.6 can be taken over all of $C^{0}(X)$ or over $W^{2,1}(X).$ Indeed, by classical Hilbert space theory the latter sup is attained by $u_{\mu}\in W^{2,1}(X)$ solving the (Laplace) equation 3.7 and hence $u_{\mu}\in\text{PSH}(X,\omega_{0})$.
@@ 3.2.5. The space $\mathcal{R}_{1,1}(X,\omega)$ of geodesics rays in $\mathcal{H}_{1,1}(X,\omega)$
The space $\mathcal{H}(X,\omega_{0})$ of Kähler potentials can be endowed with a canonical Riemannian metric, first introduced by Mabuchi, where the squared length of a tangent vector $\dot{u}$ at $u$ is defined by $\int_{X}\dot{u}^{2}\omega_{u}^{n}.$ The corresponding notion of geodesics was extended to the space
in [40]. We will denote by $\mathcal{R}(X,\omega)_{1,1}$ the corresponding space of all geodesic rays $u_{t}$ in $\mathcal{H}_{1,1}(X,\omega),$ parametrized by $t\in[0,\infty[.$
3.2.6. The Mabuchi functional $\mathcal{M}$ versus the free energy functional $F_{\beta}$
@@
We recall that the function $(\mu,\nu)\mapsto D(\mu|\nu)$ is non-negative and lower semicontinuous (lsc) on $\mathcal{P}(X)^{2}$ and $\mu\mapsto D(\mu|\nu)$ is convex on $\mathcal{P}(X).$ These properties follow directly from the following well-known Legendre duality formula [50]:
@@ 3.2.7. The Futaki character of $(X,\Delta)$ and its quantization
More generally, 3.19 holds if $\psi$ and $\phi$ have finite energy and 3.20 holds if $(dd^{c}\psi)^{n}/V$ has finite entropy. In particular, if $(X,\Delta)$ admits a KE metric, then $\text{Fut}_{(X,\Delta)}(\cdot)\equiv 0.$
@@
This is well-known, at least for smooth data, but for completeness we provide simple proofs. To prove 3.19 first note that, for any given any $g\in\mathcal{G},$
@@
This follows directly from the fact that the differential $d\mathcal{E},$ which is the one-form on 3.5 defined by formula 3.5, satisfies $g^{*}(d\mathcal{E})=d(g^{*}\mathcal{E}).$ Replacing $g$ with a one-parameter subgroup $g_{\tau}$ and differentiating with respect to $\tau$ then proves formula 3.19. Next, observe that $\text{Fut}_{(X,\Delta)}(V^{1,0})=0$ for all $V^{1,0}\in\mathfrak{g}$ iff $\int_{X}\mathcal{L}_{\text{Re}V}\phi_{0}(dd^{c}\phi_{0})^{n}=0$ for all $V^{1,0}\in\mathfrak{g}$ (by applying $\text{Fut}_{(X,\Delta)}$ to both $V^{1,0}$ and $iV^{1,0}).$ Now, using that $\mathfrak{g}$ is reductive, we can decompose $\text{Re}V=W_{1}+JW_{2}$ where $W_{i}\in\mathfrak{k}$, the Lie algebra of $\mathcal{K}$. Since $\phi_{0}$ is assumed $\mathcal{K}$–invariant $\text{Fut}_{(X,\Delta)}(V^{1,0})=\int h_{W_{2}}(\omega_{0})^{n}/V,$ which shows that $\text{Fut}_{(X,\Delta)}(\cdot)\equiv 0$ iff $\int_{X}\boldsymbol{m}(\omega_{0})^{n}/V=0,$ i.e. iff $\omega_{0}^{n}/V\in\mathcal{P}(X)_{0}.$ Finally, applying formula 3.13, we can decompose
@@
Differentiating with respect to $t$ and using 3.4 thus proves 3.20. More generally, the proof of 3.19 only used that $\phi$ and $\psi$ have finite energy. Moreover, when $(dd^{c}\psi)^{n}$ has finite entropy it is shown in [23] that $\psi$ has finite energy, showing that $D(dd^{c}\psi)^{n}|\mu_{\psi})<\infty.$ Finally, if $(X,\Delta)$ admits a Kähler–Einstein metric $\psi,$ then, by [24], $(dd^{c}\psi)^{n}$ has finite entropy and $\psi$ minimizes the Mabuchi functional on the space of metrics with finite energy. Hence, $\frac{d}{dt}\mathcal{M}\left(\exp(t\text{Re}V^{1,0})^{*}\psi\right)_{|t=0}$, showing that $\text{Fut}_{(X,\Delta)}(\cdot)\equiv 0.$
@@ 3.2.8. The moment maps $\boldsymbol{m}$ and $\boldsymbol{m}_{N}$
where $h_{W}$ denotes the Hamiltonian of a real vector field $W$ on $X$ corresponding to an element $W$ in $\mathfrak{k},$ defined by $h_{W}\coloneqq\mathcal{L}_{JW}\phi$ (defined as in formula 3.18). We will denote by $\boldsymbol{m}_{N}$ the moment map of the diagonal action of the group $\mathcal{K}$ on $X^{N},$ endowed with the symplectic form
@@ 3.3. The structure of the space $\mathcal{P}(X)_{0}$ of probability measures with vanishing moment
and since $i\partial\bar{\partial}\phi^{\tau}$ is smooth and strictly positive on $X-\Delta$ it follows that $W=0$ on $X-\Delta.$ Thus, by continuity, $W=0$ on $X,$ which is a contradiction. Hence, 3.23 holds. Finally, the inequality 3.24 is shown in essentially the same way as in the proof of [21, Lemma 9.1]. More precisely, in [21, Lemma 9.1] it is assumed that the corresponding metric $\phi^{\tau}$ is independent of the real part of $\tau$ (in our notation) and that $\phi^{\tau}$ satisfies the complex geodesic equation, that is the local function $\Phi(z,\tau)\coloneqq\phi^{\tau}(z)$ satisfies
@@
where $h_{W}$ is the Hamiltonian function corresponding to $W$ and $\omega_{0}.$ First recall that $h_{W}\coloneqq\mathcal{L}_{JW}\phi_{0},$ (see Section 3.2.8). To prove formula 3.26 note that, by 3.10, denoting by $\phi$ the potential of $\mu,$
@@
which proves 3.26, using the previous lemma and the assumption that the Futaki character of $(X,\Delta)$ vanishes.
@@
First note that $F_{-\gamma}$ is $\mathcal{K}$–invariant, since $\phi_{0}$ is $\mathcal{K}$–invariant. Moreover, by definition, $F_{-\gamma}=(1-\gamma)E+F_{-1}.$ Since $F_{-1}$ is $\mathcal{G}$–invariant, under the assumption about the vanishing of the Futaki character, the first statement of the lemma thus follows and, as a consequence, also the first identity in formula 3.27. Thus, to prove the rest of the identities 3.27, it will be enough to show that for any given $\mu$ in $\mathcal{P}(X)$ such that $F_{\gamma}(\mu)<\infty$, there exists $\nu_{j}\in\mathcal{V}(X)$ such that $F(\nu_{j})\rightarrow F(\mu).$ But this follows from [27, Lemma 3.1], implying that for any given $\mu$ such that $E(\mu)<\infty$ and $D(\mu|\mu_{0})<\infty,$ there exists a sequence $\mu_{j}$ converging towards $\mu$ in $\mathcal{P}(X)$ such that $E(\mu_{j})$ and $D(\mu_{j}|\mu_{0})$ converge towards $E(\mu)$ and $D(\mu|\mu_{0}),$ respectively. ∎
@@
To prove the first statement of the proposition it is enough to show that when changing the metric $\phi_{0}$ the corresponding pluricomplex energy $E_{\phi_{0}}$ only changes by an additive bounded term. To this end note that, by formula 3.10, given $\phi_{0}$ and $\psi_{0}$ in $\mathcal{H}(L)$ the differential $d(E_{\phi_{0}}-E_{\phi_{1}})|_{\mu}$ is represented by the function $u\coloneqq\phi_{0}-\phi_{1},$ for any $\mu$ in $\mathcal{P}(X)$ satisfying $E(\mu)<\infty.$ Hence, there exists a constant $C$ such that
@@
showing that $\left|E_{\phi_{0}}(\mu)-E_{\phi_{1}}(\mu)\right|\leq C+\left\|u\right\|_{L^{\infty}},$ as desired. Finally, the second statement follows from combining formulae 3.13 and 3.11 with proposition 3.4 and using that $F_{-1}$ is $\mathcal{G}$–invariant (Section 3.2.7). ∎
@@
for some positive constants $A$ and $B.$ Indeed, by 3.11, $E(\frac{\omega_{u}^{n}}{V})=(\mathcal{I}-\mathcal{J})(u)$, which is comparable to $\mathcal{J}(u)$ [10, Lemma 2.13] and by [48, Prop 5.5] the inequalities 3.30 do hold when $E(\frac{\omega_{u}^{n}}{V})$ is replaced by $\mathcal{J}(u).$ This proves $1\iff 2$.
Next, note that, $3\implies 2$ follows from the previous proposition. To prove that $2\implies 3$ we recall that it was pointed out in [48, Remark 3.9] that, by density, the inequality 3.29 is equivalent to the corresponding inequality on the metric completion of the metric space $\left(\mathcal{H}(X,\omega_{0}),d_{1}\right),$ which, by [46], may be identified with $\mathcal{E}^{1}(X,\omega_{0}).$ Combined with the inequalities 3.30 this shows that (after perhaps decreasing $\epsilon$ and increasing $C$) there exist $\epsilon,C>0$ such that
@@
To prove the last statement of the theorem, recall that any KE volume form $\mu$ in $\mathcal{P}(X)$ minimizes $F_{-1}$ on $\mathcal{P}(X)$ [24]. Thus, by Lemma 3.5, there exists a KE volume form $\pi_{0}(\mu)$ minimizing $F_{-1}$ on $\mathcal{P}(X)_{0}$ (but, a priori, $\pi_{0}(\mu)$ is only determined modulo $\mathcal{K}).$ To prove that there exists a unique KE volume form $\mu_{\text{KE}}$ in $\mathcal{P}(X)_{0},$ we recall that, as shown in [21, Lemma 6.3], for any given $\phi_{0}\in\mathcal{H}(-K_{(X,\Delta)})$ there exists a unique KE volume form $\mu_{-1}$ in $\mathcal{P}(X)$ minimizing $E$ on the space of KE volume forms in $\mathcal{P}(X),$ i.e. on $\mathcal{G}\mu_{-1}.$ Moreover, $\mu_{-1}$ arises as the limit 3.28. In the present case, where $\phi_{0}$ is assumed $\mathcal{K}$–invariant, it thus follows from Prop 3.4 that $\mu_{-1}\in\mathcal{P}(X)_{0}$ and conversely, any KE volume form in $\mathcal{P}(X)_{0}$ coincides with $\mu_{-1}.$ It also follows from the aforementioned uniqueness that $\mu_{\text{KE}}$ is $\mathcal{K}$–invariant. ∎